# Spectral Dataset of Young Type Ib Supernovae and their Time-evolution

### v1

@@ 4 Supernova Classification via SNID
Ensuring a robust classification of SNe in our sample as Type Ib is an essential criterion for our data set. As discussed below in detail for each SN in our sample, we classify each SN by running the spectrum closest to the V-band peak ( = 0 days) through SNID (Blondin & Tonry 2007). We also run SNID on the early spectra (i.e., ¡ 0), including the earliest spectra taken after discovery, to verify consistency and identify SNe showing uncertain or evolving early types. SNID operates by removing the continuum from the newly obtained spectrum (in order to focus on the spectral lines that define the most common SN types and to take out the effect of any reddening) and by cross-correlating it with a library of previously observed and classified SN spectra, called individual ”SN templates”. SNID provides a quality parameter called rlap to measure the correlation; generally, correlations with rlap ¿ rlapmin = 5.0 are considered a good match (Blondin & Tonry 2007). Therefore, the SNID matches reflect spectral behavior in only the absorption lines, with stronger absorption lines having higher weights, and do not take into account the shape of the continuum. For the SNID SESN template libraries, we use the most up-to-date ones based on the Modjaz Group Sample (MGS; Modjaz et al. 2014; Liu et al. 2016a, 2017; Williamson et al. 2019a, Williamson et al. 2023) to ensure a robust classification of SNe in our sample.
@@
Classifying early spectra has been difficult due to the limited availability of previously observed young SNe Ib templates in the SNID library, or in the libraries of any of the other SN identification codes, e.g., Superfit (Howell et al., 2005), GELATO (Harutyunyan et al., 2008), for that matter. For SNID, the current MGS so far includes 235 spectral templates of 22 SNe Ib with ¡ 60 days, but only 91 spectra have ¡ 0 days. Additionally, spectra may not develop strong identifiable features, which is a well-known case for the He lines (e.g. Liu et al. 2016a, Williamson et al. 2019a,Williamson et al. 2021). Therefore, while the spectra closest to the V-band maximum light of the SNe in our sample all match those of normal SNe Ib in the SNID library (since they were chosen as such to be included in this paper), some of them may have uncertain or evolving-types before maximum light, which we are uncovering in detail in this paper. For example, as discussed below for the individual SNe, the earliest spectra of SN 2023ljf, SN 2021ukt, and SN 2021njk do not strongly match any specific SN type. Additionally, SN 2022nyo, SN 2021ukt, and SN 2019odp show evolving types in their early spectra, as described in detail below. Furthermore, we compared our classification with the TNS classification attempts: Table 2 provides the TNS classification in detail for each SN in the sample. SN 2021ukt and SN 2021njk have different classifications than Type Ib on TNS. Below, we detail the discovery and classification information for each of the SNe in our sample. For our classification information, we report the SNID classification results for the earliest spectrum and the spectrum closest to the V-band maximum light, unless otherwise mentioned.
@@ 4.1 SN 2023ljf
SN 2023ljf was discovered by the Asteroid Terrestrial Impact Last Alert System (ATLAS; Tonry et al. 2018) on UT 2023-06-22 07:40:30 (MJD 60117.32) with a non-detection two days prior (Tonry et al., 2023). Pellegrino et al. (2023), as part of the GSP, obtained a classification spectrum of the object on UT 2023-06-23 09:29:05 (t = -14.5d) and used it to classify SN 2023ljf as a young SN Ib on TNS. We ran our spectrum closest to the V-band maximum light taken on UT 2023-07-03 07:23:36 (t = -4.6d) through SNID. SNID found that the best type is a normal SN Ib, with 55.7 of the templates with rlap rlapmin corresponding to this type. We also ran our earliest spectrum taken on UT 2023-06-23 09:29:05 (t = -14.5d) through SNID. SNID did not find any favored type or subtype for this spectrum. Note that Pellegrino et al. (2023) used the same spectrum to classify SN 2023ljf as a normal SN Ib on TNS. However, they do not mention the use of SNID for classification. For the subsequent spectrum taken on UT 2023-06-25 08:16:46 (t = -12.5d), SNID found the best match to be SN Ib, with 74.6 of the templates with rlap rlapmin corresponding to normal SN Ib.
@@ 4.2 SN 2022nyo: transitioning from SN IIb to SN Ib
SN 2022nyo was discovered by the Distance Less Than 40 Mpc (DLT40) survey (Tartaglia et al., 2018) on 2022-06-30 06:22:01 (MJD 59760.27) with a non-detection 5 days prior (Pearson et al., 2022b). On TNS, 2022nyo is a Type Ib due to the classification of Pearson et al. (2022a), who used a spectrum taken on UT 2022-07-01 00:39:45 (MJD 59761.03) taken by the Gemini South telescope. We ran our spectrum closest to the V-band maximum light, taken on UT 2022-07-15 10:06:09 (t = -1.5d) through SNID. SNID did not find a favored type match for this spectrum but found the best matches with normal SN Ib. However, the earliest spectrum of SN 2022nyo (t = -13.8d) shows a prominent H absorption feature around 6200 (implying an absorption velocity of 20,000 km/s, which is similar to other SNe IIb and SNe Ib at similar phases, see Fig. 2 in Liu et al. 2016b), which diminished in strength in the subsequent spectra starting at = +4.1 days. SNID confirms the H-like feature in the earliest spectrum by finding its best match with the Type IIb SN 2008cw (Modjaz et al., 2014) and 87.5 of the templates with rlap rlapmin corresponding to Type IIb. Although the feature at 6200 could be due to other lines, such as C II 6580, Ne I 6402, or Si II 6355 (e.g., Branch et al. 2002; Tominaga et al. 2005; Elmhamdi et al. 2006; Parrent et al. 2007), several studies suggest that is most likely some H in SNe Ib (e.g. Liu et al. 2016a). A qualitatively similar early-time transition from Type IIb to Type Ib has been noted in SN 2022crv (Dong et al., 2023a; Gangopadhyay et al., 2023), though a similarly detailed study of SN 2022nyo is outside the scope of this paper.
| SN Name | Our Classification | TNS Classification | MJD of TNS Classification Spectrum | Phase (days)a of TNS Classification Spectrum | Reference |
@@ 4.3 SN 2021ukt: Peculiar SN Ib with IIn-like features
SN 2021ukt was discovered via the Zwicky Transient Facility (ZTF; Graham et al. 2019) on 2021-07-31 10:34:47 (MJD 59426.44) with a non-detection 7 days prior (De, 2021). SN 2021ukt is a unique case within our sample given its peculiar evolution. TNS lists SN 2021ukt as a Type IIn due to the classification report of Hinkle (2021b), which was based on a quick-reduction spectrum at t=-14.1 days (1.9 days before our first spectrum) showing a strong blue continuum and narrow H and H emission lines. For our earliest spectrum taken on UT 2021-08-03 11:41:04 (t = -12.2d), SNID did not find any significant classification type for this spectrum - nevertheless, a narrow H that can be well-fit by a Lorenzian shape with a FWHM km/s (see below for the fit) is visible, along with its blue continuum, fulfilling the classification criterion of a SN IIn. For the spectrum on UT 2021-08-06 12:17:32 (t = -9.1d), SNID found matches to 19 SN templates with rlap rlapmin, without a strong match to any single type. Nevertheless, the beginning of the He absorption lines (at ca 15,000 km/s) are seen in the spectra. For the following spectrum taken on UT 2021-08-09 11:57:22 (t = -6.2d), SNID found the favored type to be normal SN Ib, with 60.0 of the templates with rlap rlapmin corresponding to this type. In addition, we ran our spectrum closest to the V-band maximum, taken on UT 2021-08-15 13:30:07 (t = -0.1d) through SNID. SNID found that the best-fitting type is a normal SN Ib, with 71.1 of the templates with rlap rlapmin corresponding to normal SN Ib. Thus, we classify SN 2021ukt as a Type Ib based on the classification closest to the V-band maximum and mentions its type change in Tables 1 and 2.
@@
SN 2021ukt displays a broad HeI 5876 absorption line which decreases in absorption velocity to 8000 km/s at 45.6 days, while the narrow H emission line persist throughout our spectra at varying strengths. As shown in Figure 2, the shape of the H emission line evolves from a Lorentzian profile in our earliest spectrum, at = -12.2 days with FWHM km/s to a two-component Gaussian profile in our latest spectrum at = +45.8 days with a similarly narrow component( FWHM km/s). There is also an intriguing redshifted broad H component in the spectrum at = +45.8 days centered at 2680 km/s, but its detailed interpretation is beyond the scope of this work. We obtained fits to the H emission line by performing MCMC fits (10000 steps with a 1000-step burn-in run) with emcee (Foreman-Mackey et al., 2013) and adopted the 16th and 84th percentiles of the posterior distribution as the upper and lower error bars on the FWHM.
@@
The combination of broad He absorption with a narrow H emission line at the same time is something that has never been seen up to this point for an SN of any type. While SN 2014C (Milisavljevic et al., 2015) has been known to evolve from a SN Ib to an SN IIn and develop strong narrow emission lines, there has never been a case when a SN so strongly displayed characteristics of both subtypes at the same time. A detailed paper on this SN, focusing on modeling its light curve to infer the properties of its progenitor, will be presented by Pichay et al. (in prep).
@@ 4.4 SN 2021njk
SN 2021njk was discovered by ATLAS on UT 2021-05-25 06:48:57 (MJD 59354.38) with a non-detection two days prior (Tonry et al., 2021) and classified by the GSP Hiramatsu et al. (2021). Given that the TNS classification spectrum, taken on UT 2021-05-30 09:51:19 (t = -11.3d), was obtained by LCO, this spectrum is also present in our study. SNID indicates the best match to SN 2005bf for this spectrum, in line with the TNS classification report by Hiramatsu et al. (2021). SN 2005bf is regarded as a peculiar Type Ib because it shows a light curve with two peaks and He lines with increasing velocity over time (Tominaga et al., 2005). However, for the purpose of spectroscopic classification, we list SN 2005bf as normal SN Ib. Hiramatsu et al. (2021) also found a match with SN 1999ex using SNID and Superfit. Although Hiramatsu et al. (2021) claimed SN 1999ex as peculiar Type Ib, Modjaz et al. (2014) argued that SN 1999ex is a normal Type Ib and not an intermediate Ib/c. Additionally, we also found top matches with other normal SN Ib (iPTF13bvn, SN 2001ai, SN 2009jf) via SNID. We also ran the spectrum closest to the V-band maximum date, taken on UT 2021-06-13 10:05:47 (t = 2.7d) through SNID. Although SNID did not find a favored type, the top 6 template matches are normal Ib. Thus, we classify SN 2021njk as a normal SN Ib.
@@ 4.5 SN 2021hen
SN 2021hen was discovered by the Automatic Learning for the Rapid Classification of Events (ALeRCE) broker team (Förster et al., 2021) on UT 2021-03-28 08:07:19.004 (MJD 59301.54) in ZTF data with a non-detection four days prior (Forster et al., 2021). SN 2021hen was initially classified as an SN I using a GSP-obtained spectrum by Pellegrino et al. (2021), who found SNID matches to both Type Ia and Type Ib SNe. This SN was reclassified as an SN Ib by Sollerman (2022) using a spectrum obtained 2.5 days after the initial classification spectrum. We ran the spectrum closest to the V-band maximum light, taken on UT 2021-04-06 11:54:36 (t = -1.2d) through SNID. SNID classified this spectrum as a normal SN Ib, with 37.2 of the templates with rlap rlapmin corresponding to this type. For our earliest spectrum taken on UT 2021-03-30 13:01:37 (t = -8.2d), SNID classified this spectrum as normal Ib, with 41.8 of the templates with rlap rlapmin corresponding to this type. Pellegrino et al. (2021) used the same spectrum for their initial classification on TNS, reporting SN 2021hen as Type Ia or Type Ib. But for our classification based on the same SN spectrum, SNID found the best match with SN Ib by using the updated MGS SNID template library with significantly more SESNe templates.
@@ 4.6 SN 2020hvp
SN 2020hvp was discovered by ATLAS on UT 2020-04-21 12:24:28 (MJD 58960.50) with a non-detection two days prior (Tonry et al., 2020). It was classified as an SN Ib by both Burke et al. (2020) and Dahiwale & Fremling (2020) using spectra obtained 2 and 4 days after discovery, respectively. We ran our spectrum closest to the V-band maximum, taken on UT 2020-05-06 14:08:49 (t = -0.4d) through SNID. SNID found that the best match is a normal SN Ib, with 29.3 of the templates with rlap rlapmin corresponding to this type. For the earliest spectrum taken on UT 2020-04-23 17:50:18 (t = -13.2d), SNID did not find any favored type or subtype, but the top 6 template matches are normal SNe Ib.

### v2

@@ 4 Supernova classification via SNID
Ensuring a robust classification of SNe in our sample as type Ib is an essential criterion for our dataset. As discussed below in detail for each SN in our sample, we classify each SN by running the spectrum closest to the V-band peak ( = 0 days) through SNID (Blondin & Tonry 2007). We also run SNID on the early spectra (i.e., ¡ 0), including the earliest spectrum taken after discovery, to verify consistency and identify SNe showing uncertain or evolving early types. SNID operates by removing the continuum from the newly obtained spectrum (in order to focus on the spectral lines that define the most common SN types and to take out the effect of any reddening) and by cross-correlating it with a library of previously observed and classified SN spectra, called individual ”SN templates”. SNID provides a quality parameter called rlap to measure the correlation; generally, correlations with rlap ¿ rlapmin = 5.0 are considered a good match (Blondin & Tonry 2007). Therefore, the SNID matches reflect spectral behavior in only the absorption lines, with stronger absorption lines having higher weights, and do not take into account the shape of the continuum. For the SNID SESN template libraries, we use the most up-to-date ones based on the Modjaz Group Sample (MGS; Modjaz et al. 2014; Liu et al. 2016, 2017; Williamson et al. 2019, 2023) to ensure a robust classification of SNe in our sample.
@@
Classifying early spectra has been difficult due to the limited availability of previously observed young SNe Ib templates in the SNID library, or in the libraries of any of the other SN identification codes, such as Superfit (Howell et al., 2005) and GELATO (Harutyunyan et al., 2008), for that matter. For SNID, the current MGS so far includes 235 spectral templates of 22 SNe Ib with ¡ 60 days, but only 91 spectra have ¡ 0 days. Additionally, spectra may not develop strong identifiable features, which is a well-known case for the He lines (e.g., Liu et al. 2016; Williamson et al. 2019, 2021). Therefore, while the spectra closest to the V-band maximum light of the SNe in our sample all match those of normal SNe Ib in the SNID library (since they were chosen as such to be included in this paper), some of them may have uncertain or evolving-types before maximum light, which we are uncovering in detail in this paper. For example, as discussed below for the individual SNe, the earliest spectra of SN 2023ljf, SN 2021ukt, and SN 2021njk do not strongly match any specific SN type in SNID. Additionally, SN 2022nyo, SN 2021ukt, and SN 2019odp show evolving types in their early spectra, as described in detail below. Furthermore, we compared our classification with the TNS classification attempts: Table 3 provides the TNS classification in detail for each SN in the sample. SN 2021ukt and SN 2021njk have different classifications than type Ib on TNS. Below, we detail the discovery and classification information for each of the SNe in our sample. For our classification information, we report the SNID classification results for the earliest spectrum and the spectrum closest to the V-band maximum light, unless otherwise mentioned.
The types of all the spectra in our sample are listed in Table LABEL:tab:_obs_log. However, we note that the types for the spectra in Table LABEL:tab:_obs_log are not always determinable by SNID due to the limitations of the current SNID library, particularly its incompleteness regarding early spectra (as discussed above). To enhance the reliability of our classifications, we employ complementary methods, including visual confirmation of key spectral features and the fitting of line profiles, as demonstrated in the case of SN 2021ukt (Section 4.3). We hope that our newly added SNID templates (see Section 5) will help address some of the incompleteness in the library regarding early spectra of SNe Ib.
@@ 4.1 SN 2023ljf
SN 2023ljf was discovered by the Asteroid Terrestrial Impact Last Alert System (ATLAS; Tonry et al. 2018) on UT 2023-06-22 07:40:30 (MJD 60117.32) with a non-detection two days prior (Tonry et al., 2023). Pellegrino et al. (2023), as part of the GSP, obtained a classification spectrum of the object on UT 2023-06-23 09:29:05 (t = -14.5d) and used it to classify SN 2023ljf as a young SN Ib on TNS. We ran our spectrum closest to the V-band maximum light taken on UT 2023-07-03 07:23:36 (t = 4.6d) through SNID. SNID found that the best type is a normal SN Ib, with 55.7 of the templates with rlap rlapmin corresponding to this type. We also ran our earliest spectrum taken on UT 2023-06-23 09:29:05 (t = 14.5d) through SNID. SNID did not find any favored type or subtype for this spectrum. It is important to note that Pellegrino et al. (2023) used the same spectrum to classify SN 2023ljf as a normal SN Ib on TNS. However, they do not mention the use of SNID for classification. For the subsequent spectrum taken on UT 2023-06-25 08:16:46 (t = 12.5d), SNID found the best match to be SN Ib, with 74.6 of the templates with rlap rlapmin corresponding to normal SN Ib.
@@ 4.2 SN 2022nyo: Transitioning from SN IIb to SN Ib
SN 2022nyo was discovered by the Distance Less Than 40 Mpc (DLT40) survey (Sand, 2023) on UT 2022-06-30 06:22:01 (MJD 59760.27) with a non-detection 5 days prior (Pearson et al., 2022b). On TNS, 2022nyo is a type Ib based on the classification of Pearson et al. (2022a), who used a spectrum taken on UT 2022-07-01 00:39:45 (MJD 59761.03) with the Gemini South telescope. We ran our spectrum closest to the V-band maximum light, taken on UT 2022-07-15 10:06:09 (t = 1.5d) through SNID. SNID did not find a favored type match for this spectrum but found the best matches with normal SN Ib. However, the earliest spectrum of SN 2022nyo (t = 13.8d) shows a prominent H absorption feature around 6200 (implying an absorption velocity of 20,000 km/s, which is similar to other SNe IIb and SNe Ib at similar phases, see Fig. 2 in Liu et al. 2016), which diminished in strength in the subsequent spectra starting at = 4.1 days. SNID confirms the H-like feature in the earliest spectrum by finding its best match with the type IIb SN 2008cw (Modjaz et al., 2014) and 87.5 of the templates with rlap rlapmin corresponding to type IIb. Although the feature at 6200 could be due to other lines, such as C II 6580, Ne I 6402, or Si II 6355 (e.g., Branch et al. 2002; Tominaga et al. 2005; Elmhamdi et al. 2006; Parrent et al. 2007), several studies suggest that there is most likely some H in SNe Ib (e.g., Liu et al. 2016). A qualitatively similar early-time transition from type IIb to type Ib has been noted in SN 2022crv (Dong et al., 2024; Gangopadhyay et al., 2023), though a similarly detailed study of SN 2022nyo is outside the scope of this paper.
| SN name | Our classification | TNS classification | MJD of TNS classification spectrum | Phasea (days) of TNS classification spectrum | Reference for initial classification |
@@
b Based on our spectra, SN 2022nyo transitioned from a SN IIb to a SN Ib between = 13.8 and 4.1 days (see Table LABEL:tab:_obs_log).
c Based on our spectra, SNID identified SN 2021ukt as a SN IIn at = 12.2 days, but as a SN Ib for the subsequent spectrum at = 9.1 days (see Table LABEL:tab:_obs_log). Nevertheless, we note that for many of its spectra, SN 2021ukt seems to have simultaneous SNe Ib- and SN IIn-like features, which has never been seen before—see Section 4.3.
d Based on our spectra, SN 2019odp transitioned between a SN Ic-bl and a SN Ib between = 2.7 and 1.4 days (see Table LABEL:tab:_obs_log).
@@ 4.3 SN 2021ukt: Peculiar SN Ib with IIn-like features
SN 2021ukt was discovered via the ZTF on UT 2021-07-31 10:34:47 (MJD 59426.44) with a non-detection 7 days prior (De, 2021). SN 2021ukt is a unique case within our sample given its peculiar evolution. TNS lists SN 2021ukt as a type IIn due to the classification report of Hinkle (2021b) as part of the Spectroscopic Classification of Astronomical Transients (SCAT; Tucker et al. 2022) survey, which was based on a quick-reduction spectrum at t=14.1 days (1.9 days before our first spectrum) showing a strong blue continuum and narrow H and H emission lines. For our earliest spectrum taken on UT 2021-08-03 11:41:04 (t = 12.2d), SNID did not find any significant classification type for this spectrum - nevertheless, a narrow H that is better fit by a Lorentzian shape (=1.7) vs. a Gaussian(=2.1) with a FWHM km/s (see below for the fit) is visible, along with its blue continuum, fulfilling the classification criterion of a SN IIn. For the spectrum on UT 2021-08-06 12:17:32 (t = 9.1d), SNID found matches to 19 SN templates with rlap rlapmin, without a strong match to any single type. Nevertheless, the beginning of the He absorption lines (at 15,000 km/s) are seen in the spectra. For the following spectrum taken on UT 2021-08-09 11:57:22 (t = 6.2d), SNID found the favored type to be normal SN Ib, with 60.0 of the templates with rlap rlapmin corresponding to this type. In addition, we ran our spectrum closest to the V-band maximum, taken on UT 2021-08-15 13:30:07 (t = 0.1d) through SNID. SNID found that the best-fitting type is a normal SN Ib, with 71.1 of the templates with rlap rlapmin corresponding to normal SN Ib. Thus, we classify SN 2021ukt as a type Ib based on the classification closest to the V-band maximum and mention its type change in Tables 1 and 3.
@@
SN 2021ukt displays a broad HeI 5876 absorption line which decreases in absorption velocity to 8000 km/s at 45.6 days, while the narrow H emission line persists throughout our spectra at varying strengths. As shown in Figure 2, the shape of the H emission line evolves from a Lorentzian profile in our earliest spectrum, at = 12.2 days with FWHM km/s to a two-component Gaussian profile in our latest spectrum at = +45.8 days with a similarly narrow component( FWHM km/s). There is also an intriguing redshifted broad H component in the spectrum at = +45.8 days centered at 2680 km/s, but its detailed interpretation is beyond the scope of this work. We obtained fits to the H emission line by performing MCMC fits (10000 steps with a 1000-step burn-in run) with emcee (Foreman-Mackey et al., 2013) and adopted the 16th and 84th percentiles of the posterior distribution as the upper and lower error bars on the FWHM.
@@
The combination of broad He absorption with a narrow H emission line at the same time is something that has never been seen up to this point for an SN of any type. While SN 2014C (Milisavljevic et al., 2015) has been known to evolve from a SN Ib to an SN IIn and develop strong narrow emission lines, there has never been a case when a SN so strongly displayed characteristics of both subtypes at the same time. A detailed paper on this SN, focusing on modeling its light curve to infer the properties of its progenitor, will be presented by Pichay et al. (in prep; Pichay et al. 2024).
@@ 4.4 SN 2021njk
SN 2021njk was discovered by ATLAS on UT 2021-05-25 06:48:57 (MJD 59354.38) with a non-detection two days prior (Tonry et al., 2021) and classified by the GSP Hiramatsu et al. (2021). Given that the TNS classification spectrum, taken on UT 2021-05-30 09:51:19 (t = 11.3d), was obtained by LCO, this spectrum is also present in our study. SNID indicates the best match to SN 2005bf for this spectrum, in line with the TNS classification report by Hiramatsu et al. (2021). SN 2005bf is regarded as a peculiar type Ib because it shows a light curve with two peaks and He lines with increasing velocity over time (Tominaga et al., 2005). However, for the purpose of spectroscopic classification, we list SN 2005bf as normal SN Ib. Hiramatsu et al. (2021) also found a match with SN 1999ex using SNID and Superfit. Although Hiramatsu et al. (2021) claimed SN 1999ex as peculiar type Ib, Modjaz et al. (2014) argued that SN 1999ex is a normal type Ib and not an intermediate Ib/c. Additionally, we also found top matches with other normal SN Ib (iPTF13bvn, SN 2001ai, SN 2009jf) via SNID. We also ran the spectrum closest to the V-band maximum date, taken on UT 2021-06-13 10:05:47 (t = 2.7d) through SNID. Although SNID did not find a favored type, the top 6 template matches are normal Ib. Thus, we classify SN 2021njk as a normal SN Ib.
@@ 4.5 SN 2021hen
SN 2021hen was discovered by the Automatic Learning for the Rapid Classification of Events (ALeRCE) broker team (Förster et al., 2021) on UT 2021-03-28 08:07:19.004 (MJD 59301.54) in ZTF data with a non-detection four days prior (Forster et al., 2021). SN 2021hen was initially classified as an SN I using a GSP-obtained spectrum by Pellegrino et al. (2021), who found SNID matches to both type Ia and type Ib SNe. This SN was reclassified as an SN Ib by Sollerman (2022) using a spectrum obtained 2.5 days after the initial classification spectrum. We ran the spectrum closest to the V-band maximum light, taken on UT 2021-04-06 11:54:36 (t = 1.2d) through SNID. SNID classified this spectrum as a normal SN Ib, with 37.2 of the templates with rlap rlapmin corresponding to this type. For our earliest spectrum taken on UT 2021-03-30 13:01:37 (t = 8.2d), SNID classified this spectrum as normal Ib, with 41.8 of the templates with rlap rlapmin corresponding to this type. Pellegrino et al. (2021) used the same spectrum for their initial classification on TNS, reporting SN 2021hen as type Ia or type Ib. But for our classification based on the same SN spectrum, SNID found the best match with SN Ib by using the updated MGS SNID template library with significantly more SESNe templates.
@@ 4.6 SN 2020hvp
SN 2020hvp was discovered by ATLAS on UT 2020-04-21 12:24:28 (MJD 58960.50) with a non-detection two days prior (Tonry et al., 2020). It was classified as an SN Ib by both Burke et al. (2020) and Dahiwale & Fremling (2020) using spectra obtained 2 and 4 days after discovery, respectively. We ran our spectrum closest to the V-band maximum, taken on UT 2020-05-06 14:08:49 (t = 0.4d) through SNID. SNID found that the best match is a normal SN Ib, with 29.3 of the templates with rlap rlapmin corresponding to this type. For the earliest spectrum taken on UT 2020-04-23 17:50:18 (t = 13.2d), SNID did not find any favored type or subtype, but the top 6 template matches are normal SNe Ib.